# Statistical Response of ENSO Complexity to Initial Value and Model Parameter Perturbations

### v1

@@ 4.1 Statistical response to the perturbations of initial values
The following conclusions can be made from the result using the exact formula (2) (Panel (a)). Overall, the statistical response at future time instants heavily relies on the initial state of the system. The initial perturbation has the most far-reaching impact on the subsequent evolution of the system when the initial phase is at the peak of strong EP events, where the response can remain considerable even after two years. This is unsurprising as a strong initial value can be easily amplified for a chaotic system as time goes on (Fang and Chen, 2023). The finding implies that the increased strength of El Niño events under the climate change scenario not only affects the environment of those years but also has direct subsequent impacts over a long period. In contrast, if the initial state lies at a La Niña event, the statistical response is only significant for a very short period. After that, the system follows the discharge mechanism, which is more predictable (Sharmila et al., 2023). In such a case, the difference in the initial value, namely the perturbation, is often damped quickly. Furthermore, if the system is initially in a neutral state, then the amplitudes of different state variables are all near zero. Consequently, the perturbation, which is a percentage of the initial value, is insignificant and the subsequent response becomes negligible. Qualitatively similar results are found in $T_{E}$, although the response becomes more significant for $T_{E}$ when the perturbation is imposed at the initial phases of strong EP events. Finally, the phase-locking properties are preserved in the statistical response.
@@
Figure 2 utilizes the principal coordinate direction (PCD) to demonstrate the most sensitive direction of perturbation at different lead times computed from the quad form w/ mean and variance. Note that the exact direction is given by a $6\times 1$ vector consisting of the components for the state variables $(u,h_{W},T_{C},T_{E},\tau,I)$. Yet, for the convenience of presentation, only the most significant component (in absolute value) in this 6-dimensional vector is used. This is named PCD. Later, the full eigenvector will be used to describe the most sensitive direction in the case studies. As was mentioned at the end of Section 3.3, the calculation is based on the non-dimensional system where the amplitudes of all the state variables are roughly the same. Thus, the largest component in this vector indeed reflects the dominant direction. Panels (a)–(b) show the PCD for all events across the 37-year observational period at different lead times when the statistical responses of the $T_{C}$ and $T_{E}$ are evaluated, respectively. Panels (d)–(e) summarize the schematic structures of the PCD for different events. As shown in Panel (d), perturbing $T_{C}$ always gives itself the strongest response at a short lead time, usually within three months. Yet, before the El Niño events, especially the strong EP events, thermocline depth becomes the predominant component that triggers the strongest response of $T_{C}$. A stronger thermocline depth strengthens the recharge mechanism and changes the SST patterns, affecting the EP and the CP regions. The direction along the thermocline depth is also the choice for the initial perturbation to maximize the response in the future at the interannual time scale. Next, if the perturbation is imposed when a La Niña event transits to an El Niño one, the zonal advection can play an essential role in the response at around five months lead time. This is unsurprising as the advection helps accelerate the recharge process and modify the patterns at such a time scale (An et al., 1999; Tao et al., 2023). The most complicated scenario is the multi-year La Niña events, for example, 1999-2000, where $T_{C}$ and $h_{W}$ alternate as the most sensible variable from 5 to 15 months lead time, depending on when the perturbation is imposed. In contrast, as is shown in Panel (e), when the statistical response of $T_{E}$ is considered before an EP El Niño event, perturbing $T_{E}$ gives the strongest response of itself for a short lead time. The PCD becomes $h_{W}$ as lead time increases. This simple structure is changed at the phase when an EP El Niño transits to a La Niña, where the advection again becomes important around a lead time of 5 months. For multi-year La Nina events, $T_{E}$ and $h_{W}$ alternate at different starting months as the most crucial variable to perturb that triggers the strongest response of $T_{E}$. Finally, Panel (c) shows the PCD based on the statistical response of $T_{E}$ but using the quad form w/ mean only method. The patterns in Panels (b) and (c) are similar for lead times of less than 6 months. Some differences can be seen around 7 to 9 months. The difference becomes more significant for longer ranges after 14 months, though the amplitude of the response is negligible since the initial effect almost goes away at such a long range.
@@ 4.3.2 Response to model parameter perturbations
It is worth highlighting that when computing the most sensitive perturbations with parameter perturbations, the response in the variance is much more significant than that in the mean. A strong response in the variance and higher-order statistical moments implies the increased probability of extreme events and the enhancement of the uncertainty from the model output. This indicates the fundamental difference between statistical and traditional trajectory-wise methods in studying the model response and seeking the most sensitive direction.
@@
Finally, Table 1 shows the occurrence frequency of different ENSO events in observations (first row), original system (second row), and the perturbed systems by perturbing each listed parameter by 30% (third to seventh rows). The observational period is from 1950 to 2020, totaling 71 years. When perturbing $\beta_{E}$ and $\beta_{C}$, the EP and CP El Niño events will become more frequent. Note that the number of extreme EP events is significantly increased when the wind stress or the wind amplitude is amplified, corresponding to an enhanced $\beta_{E}$. It also triggers more multi-year La Niña events. Increased occurrences of the strong El Niño and multi-year La Niña are related to the global warming scenario (Geng et al., 2023; Cai et al., 2015). Next, when perturbing $\zeta_{C}$, the feedback coefficient of $T_{E}$ in the $T_{C}$ equation, the system tends to increase the number of CP events and decrease the number of EP events, pushing more activities toward the CP region due to the strengthening of the feedback. A decrease of $\delta_{h}$, the average SSTa feedback in the thermocline depth equation, will increase the chance of multi-year El Niño events. Finally, recall that $c_{1}$ is the damping coefficient in the TC equation, representing the nonlinear parametrization of the subsurface structure as well as the discharge behavior of the SSTa in the CP region. If such a parameter is decreased, then the number of CP events will dramatically increase due to the more active response in the CP region (Zhao et al., 2021). It will also trigger more multi-year El Niño events but fewer multi-year La Niña events.
@@
Table 1: The occurrence frequency of different ENSO events in observations (first row), original system (second row), and the perturbed systems by perturbing each listed parameter by 30% (third to seventh rows). The observational period is from 1950 to 2020, totaling 71 years. A long simulation is carried out for each case. It is then divided into 30 non-overlapped segments, each being 71 years as being consistent with the length of the observations. The mean occurrence frequency numbers plus and minus the corresponding standard errors of these segments are shown.
@@ Acknowledgements.
Here we include the model parameters in the standard run (i.e., unperturbed system), and the characteristic scales appearing in the three-region multiscale stochastic model (10a)–(10f).
@@
Table 2: Parameters of the stochastic conceptual model for the ENSO complexity (3.3).

### v2

@@ 4.1 Statistical response to the perturbations of initial values
The following conclusions can be made from the result using the exact formula (2) (Panel (a)). Overall, the statistical response at future time instants heavily relies on the initial state of the system. The initial perturbation has the most far-reaching impact on the subsequent evolution of the system when the initial phase is at the peak of strong EP events, where the response can remain considerable even after two years. This is unsurprising as a strong initial value can be easily amplified for a chaotic system as time goes on (Fang and Chen, 2023). The responses from the extreme El Niños of 1982-1983 and 1997-1998 are the strongest among the different events. The initial perturbations of the other two significant El Niños, in 1987-1988 and 2015-2016, also have a long-range impact. The finding implies that the increased strength of El Niño events under the climate change scenario not only affects the environment of those years but also has direct subsequent impacts over a long period. In contrast, if the initial state lies at a La Niña event, the statistical response is only significant for a very short period. After that, the system follows the discharge mechanism, which is more predictable (Sharmila et al., 2023). In such a case, the difference in the initial value, namely the perturbation, is often damped quickly. Furthermore, if the system is initially in a neutral state, then the amplitudes of different state variables are all near zero. Consequently, the perturbation, which is a percentage of the initial value, is insignificant and the subsequent response becomes negligible. Qualitatively similar results are found in $T_{E}$, although the response becomes more significant for $T_{E}$ when the perturbation is imposed at the initial phases of strong EP events. Finally, the phase-locking properties are preserved in the statistical response.
@@
Figure 2 utilizes the principal coordinate direction (PCD) to demonstrate the most sensitive direction of perturbation at different lead times computed from the quad form w/ mean and variance. Note that the exact direction is given by a $6\times 1$ vector consisting of the components for the state variables $(u,h_{W},T_{C},T_{E},\tau,I)$. Yet, for the convenience of presentation, only the most significant component (in absolute value) in this 6-dimensional vector is used. This is named PCD. Later, the full eigenvector will be used to describe the most sensitive direction in the case studies. It will be seen in the case studies that several variables may all have non-negligible contributions at lead times ranging from 3 to 15 months. Typically, $T_{E}$ and $h_{W}$ will both contribute to the response of $T_{E}$ while $T_{C}$, $h_{W}$, $T_{E}$ and $u$ will all impact the response of $T_{C}$. The exact percentages of the contributions from each variable vary for different events. As was mentioned at the end of Section 3.3, the calculation is based on the non-dimensional system where the amplitudes of all the state variables are roughly the same. Thus, the largest component in this vector indeed reflects the dominant direction. Panels (a)–(b) show the PCD for all events across the 37-year observational period at different lead times when the statistical responses of the $T_{C}$ and $T_{E}$ are evaluated, respectively. Panels (d)–(e) summarize the schematic structures of the PCD for different events. Despite some inter-event differences, the overall patterns of each type of ENSO event are similar. As shown in Panel (d), perturbing $T_{C}$ always gives itself the strongest response at a short lead time, usually within three months. Yet, before the El Niño events, especially the strong EP events, thermocline depth becomes the predominant component that triggers the strongest response of $T_{C}$. A stronger thermocline depth strengthens the recharge mechanism and changes the SST patterns, affecting the EP and the CP regions. The direction along the thermocline depth is also the choice for the initial perturbation to maximize the response in the future at the interannual time scale. Next, if the perturbation is imposed when a La Niña event transits to an El Niño one, the zonal advection can play an essential role in the response at around five months lead time. This is unsurprising as the advection helps accelerate the recharge process and modify the patterns at such a time scale (An et al., 1999; Tao et al., 2023). The most complicated scenario is the multi-year La Niña events, for example, 1999-2000, where $T_{C}$ and $h_{W}$ alternate as the most sensible variable from 5 to 15 months lead time, depending on when the perturbation is imposed. In contrast, as is shown in Panel (e), when the statistical response of $T_{E}$ is considered before an EP El Niño event, perturbing $T_{E}$ gives the strongest response of itself for a short lead time. The PCD becomes $h_{W}$ as lead time increases. This simple structure is changed at the phase when an EP El Niño transits to a La Niña, where the advection again becomes important around a lead time of 5 months. For multi-year La Nina events, $T_{E}$ and $h_{W}$ alternate at different starting months as the most crucial variable to perturb that triggers the strongest response of $T_{E}$. Finally, Panel (c) shows the PCD based on the statistical response of $T_{E}$ but using the quad form w/ mean only method. The patterns in Panels (b) and (c) are similar for lead times of less than 6 months. Some differences can be seen around 7 to 9 months. The difference becomes more significant for longer ranges after 14 months, though the amplitude of the response is negligible since the initial effect almost goes away at such a long range.
@@ 4.3.2 Response to model parameter perturbations
It is worth highlighting that when computing the most sensitive perturbations with parameter perturbations, the response in the variance is much more significant than that in the mean. A strong response in the variance and higher-order statistical moments implies the increased probability of extreme events and the enhancement of the uncertainty from the model output. This indicates the fundamental difference between statistical and traditional trajectory-wise methods in studying the model response and seeking the most sensitive direction. It is also remarkable that the parameter perturbation will change the climatology. Notably, when the difference mainly lies in the variance and higher-order moments, the response can be insignificant when applying the trajectory-wise methods. This indicates the necessity of using the statistical response computed from the information theory.
@@
Finally, Table 1 shows the occurrence frequency of different ENSO events in observations (first row), original system (second row), and the perturbed systems by perturbing each listed parameter by 30% (third to seventh rows). The observational period is from 1950 to 2020, totaling 71 years. The results here indicate the responses of the climatology to the change of these parameters. When perturbing $\beta_{E}$ and $\beta_{C}$, the EP and CP El Niño events will become more frequent. Note that the number of extreme EP events is significantly increased when the wind stress or the wind amplitude is amplified, corresponding to an enhanced $\beta_{E}$. It also triggers more multi-year La Niña events. Increased occurrences of the strong El Niño and multi-year La Niña are related to the global warming scenario (Geng et al., 2023; Cai et al., 2015). Next, when perturbing $\zeta_{C}$, the feedback coefficient of $T_{E}$ in the $T_{C}$ equation, the system tends to increase the number of CP events and decrease the number of EP events, pushing more activities toward the CP region due to the strengthening of the feedback. A decrease of $\delta_{h}$, the average SSTa feedback in the thermocline depth equation, will increase the chance of multi-year El Niño events. Finally, recall that $c_{1}$ is the damping coefficient in the TC equation, representing the nonlinear parametrization of the subsurface structure as well as the discharge behavior of the SSTa in the CP region. If such a parameter is decreased, then the number of CP events will dramatically increase due to the more active response in the CP region (Zhao et al., 2021). It will also trigger more multi-year El Niño events but fewer multi-year La Niña events.
@@
Table 1: The occurrence frequency (i.e., the number of the events) of different ENSO events in observations (first row), original system (second row), and the perturbed systems by perturbing each listed parameter by 30% (third to seventh rows). The observational period is from 1950 to 2020, totaling 71 years. A long simulation is carried out for each case. It is then divided into 30 non-overlapped segments, each being 71 years as being consistent with the length of the observations. The mean occurrence frequency numbers (per 71 years) plus and minus the corresponding standard errors of these segments are shown.
@@ 5 Conclusion
It is worth emphasizing the different considerations when studying the two types of perturbations analyzed in this work: The initial conditions and the model parameters. The initial value perturbation will immediately impact each single ENSO event at short- and medium-range lead times. Since the role of the initial condition weakens as a function of the lead time, the long-term statistics are not affected by the initial value perturbation. In contrast, the parameter perturbation may not have a major impact at a short lead time. The response will become more significant as time evolves. It can also permanently change the climatology and is, therefore, more related to climate change. Despite this discrepancy in the goals, the response can be computed for both types of perturbations.
The study in this work highlights the response in the statistical sense. The most sensitive perturbation direction corresponds to the most influential perturbation that leads to the largest difference in the statistical forecast of the ENSO. Such difference is quantified using the information measurement, e.g., the relative entropy in Figure 1. From a broader viewpoint, both the statistical method developed here and the trajectory-wise approaches aim to find the optimal precursors that trigger the most significant change in future states. However, the statistical framework differs from many existing methods based on computing the error in the trajectories, which do not highlight the role of the uncertainty, such as the variance. Since ENSO and many other natural phenomena are chaotic and contain uncertainty, it is essential to study the difference in the optimal precursors using a single trajectory and statistics. Understanding the types of events for which the most sensitive perturbation direction will strongly depend on the variance and the higher-order moments is also extremely helpful. The most sensitive perturbation direction can also be used to improve our understanding of ENSO physics. The variables that contribute the most to these significant perturbations can be explored to discover the triggering conditions of the corresponding event and understand the gap between events in the same category but with different strengths and amplitudes.
There are a few other topics that remain as potential future work. First, as Monte Carlo methods are needed to compute the statistics, the computational cost can be increased significantly when a large number of sample points is required for operational systems. The quadratic form provides a relatively cheap practice method by computing only the leading two moments. Overall, this study shows that the resulting response using such an approximation has a similar behavior to that of using the full distribution. Yet, a more rigorous quantification of the potential errors introduced in the approximation can be an interesting topic. Second, the implication of the statistical response on the statistical forecast can be further explored. Particularly, it is interesting to study the relationship between the long-term statistical response and the potential predictability of the ENSO. Third, the statistical response can be utilized in the multi-model scenario, potentially advancing the model selection and quantifying the model error.
@@ Acknowledgements.
In Table 2 we include the model parameters in the standard run (i.e., unperturbed system), and the characteristic scales appearing in the three-region multiscale stochastic model (10a)–(10f).
@@
Table 2: Parameters of the stochastic conceptual model for the ENSO complexity (10).
Maximum entropy principle and coarse-grained statistical measurements
When studying a stochastic dynamical system we are interested in some statistical quantities which we can measure through a family of $L$ functionals that represent different statistics of the dynamics through $\mathbf{x}_{t}$, which we denote by $\mathbf{E}_{L}(\mathbf{x}_{t})=(E_{1}(\mathbf{x}_{t}),\ldots,E_{L}(\mathbf{x}_{t}))^{\mathtt{\,T}}$. At each lead time, we can extract $L$ measurements observed from the present dynamics or simulated by the model that correspond to our measurement functionals of interest, which we denote by $\overline{\mathbf{E}}_{L}=\left(\overline{E}_{1},\ldots,\overline{E}_{L}\right)^{\mathtt{\,T}}$. This measured information of the dynamics acts as a restriction, with $\overline{\mathbf{E}}_{L}$ imposing $L$ constraints that are defined through the functionals $\mathbf{E}_{L}(\mathbf{x}_{t})$, with each component defined as
$\overline{E}_{l}=\int_{\mathbf{x}_{t}}E_{l}(\mathbf{x}_{t})p^{\delta}(\mathbf{x}_{t})\mathrm{d}\mathbf{x}_{t},\ l=1,\ldots,L.$
A natural choice for $E_{l}(\mathbf{x}_{t})$ is the multivariate centralised moment of $\mathbf{x}_{t}$ of order $l=1,\ldots,L$ given by
$\mu^{\delta}_{t,l}:=\begin{cases}\int_{\mathbf{x}_{t}}\left\|\mathbf{x}_{t}\right\|p^{\delta}(\mathbf{x}_{t})\mathrm{d}\mathbf{x}_{t},&l=1\\ \int_{\mathbf{x}_{t}}\left\|\mathbf{x}_{t}-\overline{\mathbf{x}}_{t}\right\|^{l}p^{\delta}(\mathbf{x}_{t})\mathrm{d}\mathbf{x}_{t},&l\geq 2,\end{cases}$
where $\overline{\mathbf{x}}_{t}=\mathbb{E}_{p}\left[\mathbf{x}_{t}\right]$ and the expectation is taken with respect to the (true) unperturbed dynamics. As such, $E_{1}(\mathbf{x}_{t})=\left\|\mathbf{x}_{t}\right\|$ and $E_{l}(\mathbf{x}_{t})=\left\|\mathbf{x}_{t}-\overline{\mathbf{x}}_{t}\right\|^{l},l\geq 2$. In the main text we had $E_{1}(\mathbf{x}_{t})=\mathbf{x}_{t}=\mathbf{x}_{t}^{1}$ and $E_{l}(\mathbf{x}_{t})=\left(\mathbf{x}_{t}-\overline{\mathbf{x}}_{t}\right)^{2}$. This vector exponentiation can be equivalently interpreted as either the Euclidean norm raised to said power (as above), or as the product of the elements of the vector when applying that exponent element-wise. For either choice the quantities retrieved are equivalent, i.e. each measure can bound the other via a constant depending only on $l$, choice of norm, and $\mbox{Dim}(\mathbf{x}_{t})$ (Stuart and Keith, 2010).
For $\pi$ being a PDF we define its differential entropy as
$\mathcal{S}(\pi)=-\int_{\mathbf{x}}\pi(\mathbf{x})\log(\pi(\mathbf{x}))\mathrm{d}\mathbf{x}.$
By empirical information theory (Jaynes, 1957; Majda and Wang, 2006), we can then construct at each lead time the unique least biased PDF under the Maximum Entropy Principle (MEP), $p^{\delta}_{L}$, which is the PDF that maximizes the differential entropy under the constraints of the $L$ observed measurements $\overline{\mathbf{E}}_{L}$. For general $L$ constraints imposed by $\overline{\mathbf{E}}_{L}$ and defined by the functionals $\mathbf{E}_{L}(\mathbf{x}_{t})$, we have that the maximum entropy distribution is a member of the $L$-parameter exponential family of distributions (Majda and Wang, 2006), given as
$p^{\delta}_{L}(\mathbf{x}_{t})=e^{-\alpha_{0}-\bm{\alpha}_{L}\cdot\mathbf{E}_{L}(\mathbf{x}_{t})},$
where $\bm{\alpha}_{L}=\bm{\alpha}_{L}(t)=(\alpha_{1}(t),\ldots,\alpha_{L}(t))^{T}$ are the $L$ Lagrange multipliers chosen such that
$\overline{\mathbf{E}}_{L}=\int_{\mathbf{x}_{t}}\mathbf{E}_{L}(\mathbf{x}_{t})e^{-\alpha_{0}-\bm{\alpha}_{L}\cdot\mathbf{E}_{L}(\mathbf{x}_{t})}\mathrm{d}\mathbf{x}_{t},$
while $\alpha_{0}=\alpha_{0}(t)$ is determined via the normalisation condition
$e^{\alpha_{0}}=\int_{\mathbf{x}_{t}}e^{-\bm{\alpha}_{L}\cdot\mathbf{E}_{L}(\mathbf{x}_{t})}\mathrm{d}\mathbf{x}_{t}.$
Under the information theory-based framework, $\mathcal{P}(p^{\delta}(\mathbf{x}_{t}),p(\mathbf{x}_{t}))$ precisely quantifies the statistical response of a turbulent system due to the effects of a perturbation imposed on the initial state, internal fluctuations, or external forcings driving the dynamics. This quantification implicitly assumes the full statistical knowledge of the unperturbed dynamics and the availability of the perturbed densities. Such an assumption is unfeasible both practically and theoretically, thus making the need to use coarse-grained measurements, either observed or model-simulated ones, in conjunction with the MEP, necessary to construct the least-biased PDF under these constraints. For the unperturbed dynamics we can use $p_{L}(\mathbf{x}_{t})$, constructed by the MEP and the observational data from the unperturbed dynamics, as a surrogate, while for $p^{\delta}(\mathbf{x}_{t})$ we can use the density constructed by the MEP and $L^{\prime}$ model-averaged measurements from a perturbed simulation instead, where $L^{\prime}\leq L$, which we denote by $p^{M,\delta}_{L^{\prime}}(\mathbf{x}_{t})$. This is because, in practice, the model density is determined by no more information than that available in the observations. In this case, we can then see that the hurdle of insufficient knowledge can be circumvented to produce the following relation
$\mathcal{P}\left(p^{\delta}(\mathbf{x}_{t}),p_{L}(\mathbf{x}_{t})\right)=\left(\mathcal{S}\left(p^{M,\delta}_{L^{\prime}}\right)-\mathcal{S}\left(p^{\delta}\right)\right)+\mathcal{P}\left(p^{M,\delta}_{L^{\prime}}(\mathbf{x}_{t}),p_{L}(\mathbf{x}_{t})\right).$
[…]
We end this exposition by proving (7). Since the observed statistics $\overline{\mathbf{E}}_{L}$ determine the perturbed MEP density $p^{\delta}_{L}$, with the Lagrange multipliers $\bm{\alpha}_{L}$ and normalisation constant $\alpha_{0}$, by differentiating the MEP density with respect to $\bm{\delta}$ we then have that
$\cfrac{\left(\bm{\delta}\cdot\nabla_{\bm{\delta}}p_{L}^{\delta}(\mathbf{x}_{t})\right)^{2}}{p_{L}^{\delta}(\mathbf{x}_{t})}=\left((\bm{\delta}\cdot\nabla_{\bm{\delta}})\left(\alpha_{0}+\bm{\alpha}_{L}\cdot\mathbf{E}_{L}(\mathbf{x}_{t})\right)\right)^{2}p_{L}^{\delta}(\mathbf{x}_{t}),$
where we use the fact that the family of measurement functionals in question does not depend on the perturbation (under some regularity assumptions). The gradients above are calculated at the unperturbed state of $\bm{\delta}=\bm{0}$. Further differentiating $\alpha_{0}$ with respect to $\bm{\delta}$ and using a bit of vector calculus yields
$\nabla_{\bm{\delta}}\alpha_{0}$ $=e^{-\alpha_{0}}\int_{\mathbf{x}_{t}}\nabla_{\bm{\delta}}\left(e^{-\bm{\alpha}_{L}\cdot\mathbf{E}_{L}(\mathbf{x}_{t})}\right)\mathrm{d}\mathbf{x}_{t}$
$=\int_{\mathbf{x}_{t}}\nabla_{\bm{\delta}}\left(-\bm{\alpha}_{L}\cdot\mathbf{E}_{L}(\mathbf{x}_{t})\right)e^{-\alpha_{0}-\bm{\alpha}_{L}\cdot\mathbf{E}_{L}(\mathbf{x}_{t})}\mathrm{d}\mathbf{x}_{t}$
$=-\big{(}\nabla_{\bm{\delta}}\bm{\alpha}_{L}\big{)}^{\mathtt{T}}\int_{\mathbf{x}_{t}}\mathbf{E}_{L}(\mathbf{x}_{t})p^{\delta}_{L}(\mathbf{x}_{t})\mathrm{d}\mathbf{x}_{t}$
$=-\big{(}\nabla_{\bm{\delta}}\bm{\alpha}_{L}\big{)}^{\mathtt{T}}\,\overline{\mathbf{E}}_{L},$
As for the $L\times N$ Jacobian matrix $\nabla_{\bm{\delta}}\bm{\alpha}_{L}$, with standard element $\left(\nabla_{\delta}\bm{\alpha}_{L}\right)^{j=1,\ldots,N}_{l=1,\ldots,L}=\frac{\partial\alpha_{l}}{\partial\delta_{j}}$, we can get an expression for it implicitly by differentiating $\overline{\mathbf{E}}_{L}$ with respect to $\bm{\delta}$, to end up with
$\nabla_{\bm{\delta}}\overline{\mathbf{E}}_{L}$ $=\nabla_{\bm{\delta}}\left(\int_{\mathbf{x}_{t}}\mathbf{E}_{L}(\mathbf{x}_{t})p_{L}^{\delta}(\mathbf{x}_{t})\mathrm{d}\mathbf{x}_{t}\right)$
$=\int_{\mathbf{x}_{t}}\mathbf{E}_{L}(\mathbf{x}_{t})\left(-\nabla_{\bm{\delta}}\alpha_{0}-\nabla_{\bm{\delta}}\left(\bm{\alpha}_{L}\cdot\mathbf{E}_{L}(\mathbf{x}_{t})\right)\right)^{\mathtt{T}}p_{L}^{\delta}(\mathbf{x}_{t})\mathrm{d}\mathbf{x}_{t}$
$=-\int_{\mathbf{x}_{t}}\mathbf{E}_{L}(\mathbf{x}_{t})\left(-\big{(}\nabla_{\bm{\delta}}\bm{\alpha}_{L}\big{)}^{\mathtt{T}}\,\overline{\mathbf{E}}_{L}+\big{(}\nabla_{\bm{\delta}}\bm{\alpha}_{L}\big{)}^{\mathtt{T}}\mathbf{E}_{L}(\mathbf{x}_{t})\right)^{\mathtt{T}}p_{L}^{\delta}(\mathbf{x}_{t})\mathrm{d}\mathbf{x}_{t}$
$=-\left[\int_{\mathbf{x}_{t}}\mathbf{E}_{L}(\mathbf{x}_{t})\left(\mathbf{E}_{L}(\mathbf{x}_{t})-\overline{\mathbf{E}}_{L}\right)^{\mathtt{T}}p_{L}^{\delta}(\mathbf{x}_{t})\mathrm{d}\mathbf{x}_{t}\right]\nabla_{\bm{\delta}}\bm{\alpha}_{L}$
$=-\left[\int_{\mathbf{x}_{t}}\left(\mathbf{E}_{L}(\mathbf{x}_{t})-\overline{\mathbf{E}}_{L}\right)\left(\mathbf{E}_{L}(\mathbf{x}_{t})-\overline{\mathbf{E}}_{L}\right)^{\mathtt{T}}p_{L}^{\delta}(\mathbf{x}_{t})\mathrm{d}\mathbf{x}_{t}\right]\nabla_{\bm{\delta}}\bm{\alpha}_{L}$
$\equiv-\mathcal{C}\nabla_{\bm{\delta}}\bm{\alpha}_{L}.$
If we now assume invertibility of the (climate) correlation matrix $\mathcal{C}$ we have,
$\nabla_{\bm{\delta}}\bm{\alpha}_{L}=-\mathcal{C}^{-1}\nabla_{\bm{\delta}}\overline{\mathbf{E}}_{L}.$
Plugging in the expressions above for $\nabla_{\bm{\delta}}\alpha_{0}$ and $\nabla_{\bm{\delta}}\bm{\alpha}_{L}$ into $\left((\bm{\delta}\cdot\nabla_{\bm{\delta}})\left(\alpha_{0}+\bm{\alpha}_{L}\cdot\mathbf{E}_{L}(\mathbf{x}_{t})\right)\right)^{2}p_{L}^{\delta}(\mathbf{x}_{t})$ and using the definition of the Fisher information matrix and symmetricity of the correlation matrix $\mathcal{C}$ and its inverse, we then have the following series of equalities, where all gradients are calculated at the unperturbed state of $\bm{\delta}=\bm{0}$,
$\bm{\delta}\cdot I(p(\mathbf{x}_{t}))\bm{\delta}$ $=\bm{\delta}\cdot\left(\int_{\mathbf{x}_{t}}\left(-\nabla_{\bm{\delta}}\alpha_{0}-\nabla_{\bm{\delta}}\left(\bm{\alpha}_{L}\cdot\mathbf{E}_{L}(\mathbf{x}_{t})\right)\right)\left(-\nabla_{\bm{\delta}}\alpha_{0}-\nabla_{\bm{\delta}}\left(\bm{\alpha}_{L}\cdot\mathbf{E}_{L}(\mathbf{x}_{t})\right)\right)^{\mathtt{T}}p_{L}^{\delta}(\mathbf{x}_{t})\mathrm{d}\mathbf{x}_{t}\right)\bm{\delta}$
$=\bm{\delta}\cdot\left(\int_{\mathbf{x}_{t}}\big{(}\nabla_{\bm{\delta}}\overline{\mathbf{E}}_{L}\big{)}^{\mathtt{T}}\mathcal{C}^{-1}\left(\mathbf{E}_{L}(\mathbf{x}_{t})-\overline{\mathbf{E}}_{L}\right)\left(\mathbf{E}_{L}(\mathbf{x}_{t})-\overline{\mathbf{E}}_{L}\right)^{\mathtt{T}}\mathcal{C}^{-1}\nabla_{\bm{\delta}}\overline{\mathbf{E}}_{L}p_{L}^{\delta}(\mathbf{x}_{t})\mathrm{d}\mathbf{x}_{t}\right)\bm{\delta}$
$=\bm{\delta}\cdot\left(\big{(}\nabla_{\bm{\delta}}\overline{\mathbf{E}}_{L}\big{)}^{\mathtt{T}}\mathcal{C}^{-1}\nabla_{\bm{\delta}}\overline{\mathbf{E}}_{L}\right)\bm{\delta},$
which proves (7). The result and proof of this theorem generalise mutatis mutandis to the case of models with complex-valued processes and complex-valued statistical quantities of interest, by replacing the regular transpose with the conjugate transpose, since the climate correlation matrix (and by extension its inverse) is, in general, Hermitian.